Consider the function f(x)=7(x−4)^2/3

f(x)=7(x−4)2/3. For this function there are two important intervals: (−∞,A)

(−∞,A) and (A,∞)

(A,∞) where A

A is a critical number.


1
Expert's answer
2021-11-22T12:37:45-0500

"f(x)=7(x-4)^{2\/3}"

Domain: "(-\infin, \infin)"


"f'(x)=(7(x-4)^{2/3})' =7(\dfrac{2}{3})(x-4)^{-1/3}=\dfrac{14}{3}(x-4)^{-1/3}"

"=\dfrac{14}{3}(x-4)^{-1/3}"

Find the critical number(s)

"f'(x)" does not exist, if "x=4."

The critical number is "4."

If "x\in (-\infin, 4), f'(x)<0, f(x)" decreases.

If "x\in (4,\infin), f'(x)>0, f(x)" increases.